OpenStudy (anonymous):

Simplify (x + 3)(x2 + 4x – 2)

OpenStudy (ajprincess):

(x+3)(x^2+4x-2) =x(x^2+4x-2)+3(x^2+4x-2) =x*x^2+x*4x-2*x+3*x^2+3*4x-3*2 see if u can do the rest.

OpenStudy (anonymous):

okay thank you

OpenStudy (ajprincess):

yw.:)

OpenStudy (anonymous):

The rule with parenthesis is always just to multiply each thing in the first by each thing in the second.